# Spike-timing-dependent Hebbian learning as noisy gradient descent

### v2

@@ A.4 On the connection to entropic mirror descent
$=\frac{p_{i}(k)\exp(\alpha(p_{i}(k)-\|\mathbf{p}(k)\|^{2}))}{\sum_{j=1}^{d}p_{j}(k)\exp(\alpha(p_{j}(k)-\|\mathbf{p}(k)\|^{2}))}$
$\approx\frac{p_{i}(k)(1+\alpha(\nabla L(\mathbf{p}(k)))_{i})}{\sum_{j=1}^{d}p_{j}(k)(1+\alpha(\nabla L(\mathbf{p}(k)))_{j})}$
@@ A.6 Additional material for the extension to time-inhomogeneous intensities
Let $T>0$ be the time horizon and assume that the time-inhomogeneous intensities $\boldsymbol{\lambda}_{\alpha}$ change on the same scale as the weights and are given by $\boldsymbol{\lambda}_{\alpha}(k)=\boldsymbol{\lambda}(k\alpha)$, $k=0,1,\dots,\lfloor T/\alpha\rfloor$, where $\boldsymbol{\lambda}\in C_{b}^{2}([0,T])$ is a universal, twice differentiable function with bounded derivatives and $\boldsymbol{\lambda}(t)\geq\lambda_{\min}>0$, $t\in[0,T]$, componentwise. For vectors $\mathbf{w},\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}$ that are of the same length, and $\Delta:=\boldsymbol{\lambda}^{\prime}-\boldsymbol{\lambda},$ we have
@@
Sending $\alpha\to 0$ and using that $\Delta\boldsymbol{\lambda}_{\alpha}(k)/\alpha-\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\boldsymbol{\lambda}(k\alpha)\to 0$ as $\alpha\to 0$ we recognize Eq. (36) as an Euler-type scheme for the ODE
@@
$=\mathbf{p}(t)\odot\left(\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\log\big(\boldsymbol{\lambda}(t)\big)+\mathbf{p}(t)-\mathbf{p}(t)^{\top}\left(\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\log\big(\boldsymbol{\lambda}(t)\big)+\mathbf{p}(t)\right)\mathbf{1}\right),$

### v3

@@ A.4 On the connection to entropic mirror descent
$\approx\frac{p_{i}(k)(1+\alpha p_{i}(k))}{\sum_{j=1}^{d}p_{j}(k)(1+\alpha p_{j}(k))}$
@@ A.6 Additional material for the extension to time-inhomogeneous intensities
Let $T>0$ be the time horizon and assume that the time-inhomogeneous intensities $\boldsymbol{\lambda}_{\alpha}$ change on the same scale as the weights and are given by $\boldsymbol{\lambda}_{\alpha}(t)=\boldsymbol{\lambda}(t\alpha)$, $t\in[0,T/\alpha]$, where $\boldsymbol{\lambda}\in C_{b}^{2}([0,T])$ is a universal, twice differentiable function with bounded derivatives and $\boldsymbol{\lambda}(t)\geq\lambda_{\min}>0$, $t\in[0,T]$, componentwise. For vectors $\mathbf{w},\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}$ that are of the same length, and $\Delta:=\boldsymbol{\lambda}^{\prime}-\boldsymbol{\lambda},$ we have
@@
Sending $\alpha\to 0$ and using that $\Delta\boldsymbol{\lambda}_{\alpha}(t/\alpha)/\alpha\to\frac{\mathrm{d}}{\mathop{}\!\mathrm{d}t}\boldsymbol{\lambda}(t)$ as $\alpha\to 0$ we recognize Eq. (36) as an Euler-type scheme for the ODE
@@
$=\mathbf{p}(t)\odot\left(\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\log\big(\boldsymbol{\lambda}(t)\big)+\mathbf{p}(t)-\mathbf{p}(t)^{\top}\left(\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\log\big(\boldsymbol{\lambda}(t)\big)+\mathbf{p}(t)\right)\mathbf{1}\right),\quad t\in[0,T],$